OpenStudy (anonymous):

Help! 1. A company makes video games. The price of a video game is modeled by the function p(x) = 30x + 2, where x is the number of years since the company started producing games. The number of video games they sell is modeled by the function s(x) = 500x + 250. To find the total revenue from their video games, the company should use what operation on the polynomials? Addition Subtraction Multiplication It cannot be determined 2. Which of the following represents the area of a rectangle whose length is 3x + 5 and whose width is x - 2? 3x^2-x-10 3x^2-10 3x^2-x+10 3x^2-11x-10

OpenStudy (anonymous):

p(x) - s(x)=0, subtract

OpenStudy (anonymous):

x=27/470… thats what i got

OpenStudy (anonymous):

@dpasingh

OpenStudy (anonymous):

Pls wait..

OpenStudy (anonymous):

I'm sorry the functions has to be mulitplied

OpenStudy (anonymous):

ok, I'm sorry i just need someone to help explain to me how to do it.

OpenStudy (anonymous):

wait so is it subtraction or multiplication?

OpenStudy (anonymous):

total revenue = p(x) * s(x)

OpenStudy (anonymous):

oh, ok. what about the other one?

OpenStudy (anonymous):

Multiply length by width Area = (3x + 5)( x - 2) \[\huge =3x(x-2) +5(x-2) = 3x^2 -6x + 5x-10 \] \[\huge =3x^2 -x -10 \] Hence option 1 is the correct answer.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

Also i have one more question could you help me with it?

OpenStudy (anonymous):

Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x).

OpenStudy (anonymous):

how do i do this one?